OpenStudy (anonymous):

Need help in Equation.... Solve: 6y + 15 = -6 + 3y y=?

OpenStudy (anonymous):

okay we need to bring all the y terms on one side and all the constants on the other.

OpenStudy (anonymous):

can you do that step?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

that's fine. let's do this one step at a time.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we have 6y + 15 = -6 + 3y lets subtract 3y on both sides 6y-3y +15=-6+3y-3y what will this simplify to?

OpenStudy (anonymous):

i dont really know this, im bad in eqations.

OpenStudy (anonymous):

okay 6y-3y are like terms as the both have y so because of this we just subtract the numbers to get the answer. 6y-3y=3y what about 3y-3y, what do you think it will equal?

OpenStudy (anonymous):

0 ?

OpenStudy (anonymous):

0y

OpenStudy (anonymous):

because 0*anything=0 we can just write 0 perfect, so in our equation we now have 3y+15=-6+0 3y+15=-6 now we want to get the 15 to the other side, so i'm going to subtract both sides by 15 3y+15-15=-6-15 what will I get?

OpenStudy (anonymous):

-7 ?

OpenStudy (anonymous):

close, remember there is a negative in front of the 6 so -6-15=-21

OpenStudy (anonymous):

ohh i subtract it , i get it

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

3y=-21 to get rid of the 3 we need to divide both sides by 3 and then we will have an answer for y \[\frac{3y}{3}=\frac{-21}{3}\] what is y equal to?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

is it negative 7 ? or just positive ?

OpenStudy (anonymous):

don't forget about the negatives so it will be -7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you can always check by putting your value for y into the original problem